78.0 g of nonionizing C6H12O6 are dissolved in 300 g of H2O.

What is the change in boiling temperature in 0C, (ΔTb)?




1
Expert's answer
2022-05-13T05:33:03-0400

∆T = imKb

∆T = change in temperature

i = van't Hoff factor = 1 for glucose since it does not ionize or dissociate. It is a single particle.

Kb = 0.512 0C/m

m = molality = moles solut/kg solvent = (78.0 g/180 g/mol) / 0.3 kg = 1.44 m

∆T = 1 * 1.44 m * 0.512 0C/m = 0.74 0C